### v1

# Cross-Domain Transfer and Few-Shot Learning for Personal Identifiable Information Recognition

Abstract
Accurate recognition of personally identifiable information (PII) is central to automated text anonymization. This paper investigates the effectiveness of cross-domain model transfer, multi-domain data fusion, and sample-efficient learning for PII recognition. Using annotated corpora from healthcare (I2B2), legal (TAB), and biography (Wikipedia), we evaluate models across four dimensions: in-domain performance, cross-domain transferability, fusion, and few-shot learning. Results show legal-domain data transfers well to biographical texts, while medical domains resist incoming transfer. Fusion benefits are domain-specific, and high-quality recognition is achievable with only 10% of training data in low-specialization domains.
Keywords:
Personally Identifiable Information (PII) Named Entity Recognition (NER) Cross-Domain Transfer Multi-Domain Fusion Few-Shot Learning.
11footnotetext: *Corresponding author
1 Introduction
Privacy protection has become a cornerstone of modern data governance, with regulations like the European Union’s General Data Protection Regulation (GDPR) [4] establishing strict requirements for handling personally identifiable information (PII). As organizations increasingly rely on text-based data for research, analytics, and machine learning, the need for effective text anonymization has grown correspondingly urgent. Traditional manual anonymization approaches are costly, inefficient, and prone to human error [2], driving the development of automated PII recognition systems using Natural Language Processing (NLP) techniques [9] [13] [3].
However, three critical gaps remain unaddressed. First, the cross-domain transferability of PII recognition models remains largely uncharacterized [10]. Second, organizations often possess data from multiple domains, yet the potential benefits or drawbacks of combining heterogeneous training data for PII recognition remain unclear. Third, many practical deployment scenarios involve limited training data, but minimum data requirements for effective privacy protection have not been systematically evaluated.
To address these gaps, this paper conducts a comprehensive empirical study across healthcare (I2B2), legal (TAB), and biographical (Wikipedia) domains. Our systematic evaluation across 231 experimental configurations yields the following contributions:
  • Cross-Domain Transfer Characterization: We systematically analyze transfer performance across six domain pairs, revealing domain complexity hierarchies and identifying which domains serve as effective transfer sources.
  • Multi-Domain Fusion Analysis: We demonstrate that fusion benefits vary dramatically by target domain, providing domain-specific strategies for optimal performance.
  • Sample Efficiency Validation: We establish that effective PII recognition is achievable with significantly reduced training data, revealing non-monotonic relationships between data size and performance.
These findings challenge conventional assumptions about data requirements. We provide our codes at https://github.com/George-SGY/multi-domain-pii-recognition.
2 Related Research
Text anonymization research has evolved from rule-based approaches to sophisticated neural models, with most work focusing on domain-specific solutions rather than cross-domain applications that motivate our investigation.
Domain-Specific PII Recognition: The medical domain has served as the primary testbed for automated PII recognition, benefiting from well-established datasets like I2B2 [14] and achieving remarkable performance with transformer-based approaches such as SciBERT [15] and BioBERT [1] reaching F1 scores exceeding 0.98 [7]. Legal domain research has been more limited, with Pilán et al. [12] introducing the Text Anonymization Benchmark (TAB) for GDPR-compliant legal document anonymization. Other domains have seen sporadic development, including email communication datasets [6] [8] and dialogue corpora [11] [5].
Cross-Domain and Multi-Domain Approaches: Despite the proliferation of domain-specific datasets and models, systematic cross-domain evaluation remains scarce. Papadopoulou et al. [10] made a notable contribution with a Wikipedia-based biography corpus and distant supervision methods, but their work primarily focused on reducing annotation requirements rather than systematic transfer analysis. The broader NLP literature has extensively studied domain adaptation for tasks such as sentiment analysis and named entity recognition [2], establishing transfer learning as effective for leveraging knowledge across domains. However, these insights have not been systematically applied to PII recognition, where domain-specific privacy requirements and entity distributions may pose unique challenges.
3 Methodology
This section presents our comprehensive experimental framework for investigating cross-domain PII recognition. We systematically evaluate model performance across multiple domains, training paradigms, and data availability scenarios to address the four key research questions outlined in the introduction. Overall research framework is presented in Figure 1.
Figure 1: Research Framework
3.1 Datasets and Data Preparation
We selected three publicly available, human-annotated datasets representing distinct domains with varying structural characteristics and annotation schemes:
  • I2B2 (Medical Domain): 1,304 clinical narratives with 28,867 entity mentions across 23 fine-grained medical entity types. Contains 790 training and 514 test documents (average: 925 tokens).
  • TAB (Legal Domain): 1,268 legal documents with 155,006 entity mentions across 8 semantic types (PERSON, DATETIME, ORG, LOC, DEM, MISC, CODE, QUANTITY). Contains 1,014 training, 127 validation, and 127 test documents (average: 1,442 tokens).
  • Wikipedia (Biographical Domain): 553 Wikipedia biography entries with 10,714 entity mentions across 7 semantic types. Contains 453 training and 100 test documents (average: 96 tokens).
All datasets distinguish between direct identifiers (entities that uniquely identify individuals) and quasi-identifiers (contextual attributes that can reveal identity when combined). Text tokenization was performed using SpaCy with consistent handling across domains, and entity spans were converted to BIO tagging format.
3.2 Model Architectures
We benchmark four distinct model categories, each representing different approaches to PII recognition:
  • Rule-based Systems: Microsoft Presidio applying regex patterns and heuristic-based entity detection
  • Pretrained NER Models: RoBERTa-large, XLM-RoBERTa, and Flair NER models pretrained on OntoNotes 5.0, applied zero-shot without domain-specific fine-tuning
  • Domain-Tuned NER Models: The same transformer architectures (RoBERTa, Longformer, Flair) fine-tuned on domain-specific PII data. Longformer handles longer documents with its extended attention mechanism (4096 vs. 512 tokens).
  • Large Language Models (LLMs): GPT-3.5-turbo evaluated under 2-shot in-context learning.
Based on preliminary experiments, we selected Longformer as our primary model for cross-domain experiments due to its consistent performance and ability to handle varying document lengths.
3.3 Experimental Framework and Evaluation
3.3.1 Experimental Design
We conduct four complementary experiments :
  • In-Domain Evaluation: Models trained and tested within the same domain, establishing performance upper bounds and domain-specific baselines.
  • Cross-Domain Transfer: Models trained on one domain’s training set and evaluated on other domains’ test sets, assessing zero-shot generalization capabilities across the 3 by 3 domain matrix.
  • Multi-Domain Fusion: Models trained on merged corpora from multiple domains (e.g., TAB+Wikipedia, TAB+I2B2, all three domains) and evaluated on individual domain test sets, investigating whether diverse training data improves domain-specific performance.
  • Few-Shot Learning: Systematic reduction of training data volume (50, 100, 200, 300, 500+ samples) to evaluate model robustness under limited supervision and validate practical deployment scenarios.
3.3.2 Evaluation Metrics
We use token-level metrics (precision, recall, F1), mention-level metrics (exact span matching), privacy-focused metrics (direct/quasi recall, weighted precision), and entity-type analysis across semantic categories. All experiments are conducted across seven independent runs with different random seeds, reporting median scores for robust performance estimates.
4 Results and Analysis
This section presents our experimental findings across four dimensions: in-domain performance, cross-domain transferability, multi-domain data fusion, and few-shot learning feasibility.
4.1 Baseline Performance Analysis
Table 1 presents consolidated results across all model categories and domains. Domain-tuned transformer models (Longformer, RoBERTa, Flair) consistently outperform other approaches, achieving F1 scores of 0.85-0.97. Longformer demonstrates the most balanced performance with high recall while maintaining reasonable precision. Rule-based approaches (Presidio) show the weakest performance (F1: 0.53-0.65), particularly struggling with recall for direct identifiers. General NER models exhibit high recall (0.82-0.91) but suffer from low precision (0.32-0.80), resulting in excessive text masking. GPT-3.5-turbo demonstrates moderate performance (F1: 0.63-0.87) with superior precision compared to zero-shot NER models.
The medical domain (I2B2) shows the highest performance ceiling with domain-tuned models achieving F1 scores up to 0.97, while the biographical domain (Wikipedia) presents challenges in distinguishing context-dependent PII despite shorter documents. Based on these results, we select Longformer for subsequent cross-domain experiments.
Table 1: Cross-Domain Model Performance Summary. Results show F1/Recall/Precision for token-level evaluation across three domains. Best results per category are in bold.
Model Category Model TAB Wikipedia I2B2 Average F1 Rank
F1/Rec/Prec F1/Rec/Prec F1/Rec/Prec
Rule-based Presidio 0.649/0.744/0.576 0.642/0.558/0.757 0.573/0.779/0.453 0.621 10
Pretrained NER RoBERTa 0.563/0.886/0.413 0.833/0.870/0.799 0.466/0.838/0.323 0.621 10
XLM-RoBERTa 0.574/0.906/0.420 0.821/0.846/0.798 0.512/0.723/0.396 0.636 9
Flair 0.560/0.913/0.404 0.824/0.866/0.785 0.439/0.633/0.336 0.608 11
Flair-fast 0.562/0.910/0.407 0.817/0.858/0.780 0.416/0.589/0.321 0.598 12
Domain-Tuned (PII only) Longformer 0.859/0.937/0.793 0.828/0.986/0.713 0.972/0.994/0.950 0.886 1
RoBERTa 0.857/0.934/0.791 0.834/0.982/0.723 0.959/0.937/0.982 0.883 2
Flair 0.873/0.857/0.889 0.884/0.875/0.893 0.941/0.904/0.982 0.866 5
NER+ Fine-Tuned RoBERTa 0.826/0.943/0.756 0.841/0.980/0.737 0.962/0.987/0.939 0.876 4
XLM-RoBERTa 0.852/0.947/0.784 0.824/0.978/0.712 0.965/0.989/0.945 0.880 3
Flair 0.869/0.856/0.882 0.877/0.866/0.888 0.948/0.913/0.985 0.865 6
Flair-fast 0.865/0.833/0.899 0.874/0.861/0.888 0.941/0.901/0.985 0.860 7
LLM GPT-3.5-turbo 0.738/0.728/0.748 0.868/0.907/0.833 0.630/0.765/0.535 0.745 8
  • Results averaged across 7 random seeds.
  • ERdi (Direct Identifier Recall) and ERqi (Quasi Identifier Recall) metrics available in supplementary tables.
  • Domain-tuned models consistently outperform other categories, with Longformer achieving the best overall performance.
4.2 Cross-Domain Transfer Analysis
Table 2 summarizes transfer performance across all domain pairs.
TAB $\rightarrow$ Other Domains: Legal domain training demonstrates excellent transferability to biographical content (Wikipedia F1=0.847, +1.9% improvement) while showing severe degradation for medical domains (I2B2 F1=0.264, -70.8% drop).
I2B2 $\rightarrow$ Other Domains: Medical training shows surprisingly effective transfer, maintaining baseline performance on Wikipedia (F1=0.828, 0% change) and achieving reasonable performance on TAB (F1=0.738, -12.1% drop).
Wikipedia $\rightarrow$ Other Domains: Biographical training exhibits variable performance, achieving moderate results on TAB (F1=0.421, -43.8% drop) but failing dramatically on medical data (I2B2 F1=0.187, -78.5% drop).
Key Insights: (1) Transfer target hierarchy: I2B2 is most difficult to transfer into, while Wikipedia is most receptive; (2) TAB provides the most effective transfer supervision for biographical domains; (3) Medical domains resist incoming transfer but provide surprisingly effective outgoing transfer.
Table 2: Cross-Domain Transfer Performance using Longformer (Token F1 scores). Diagonal entries show in-domain baselines. Best cross-domain transfers per target are in bold.
[…]
TAB 0.859 0.847 0.264
Wikipedia 0.421 0.828 0.187
I2B2 0.738 0.828 0.972
Performance Change -32.5% +1.15% -76.8%
Transfer Direction F1 Recall Precision vs Baseline
Successful Transfers (F1 > 0.7):
TAB $\rightarrow$ Wikipedia 0.847 0.926 0.781 +1.9%
I2B2 $\rightarrow$ Wikipedia 0.828 0.803 0.825 0%
I2B2 $\rightarrow$ TAB 0.738 0.840 0.658 -12.1%
Challenging Transfers (F1 < 0.5):
Wikipedia $\rightarrow$ TAB 0.421 0.991 0.267 -43.8%
TAB $\rightarrow$ I2B2 0.264 0.990 0.152 -70.8%
Wikipedia $\rightarrow$ I2B2 0.187 0.999 0.103 -78.5%
  • Results averaged across 7 random seeds using Longformer architecture.
  • Gray cells show in-domain baselines. Performance drops show average degradation when transferring into each target domain.
4.3 Multi-Domain Fusion Effects
Table 3 presents fusion results across all domain combinations.
TAB Domain consistently improves from external data fusion, with the three-domain fusion (TWI) achieving the best recall (0.95) with modest precision decrease (-0.7% F1 overall).
Wikipedia Domain shows nuanced responses: adding I2B2 data enhances precision with minimal recall impact (+1.7% F1), while the optimal strategy depends on precision-recall priorities.
I2B2 Domain fusion consistently harms performance across all metrics (-0.2% to -0.5% F1), reinforcing the medical domain’s high specialization and sensitivity to training data noise.
Strategy Implications: Legal documents benefit from any additional training data, biographical texts require careful source selection, and medical records are best served by domain-specific training only.
Table 3: Multi-Domain Data Fusion Effects on Target Domain Performance. Results show F1/Recall/Precision for token-level evaluation using Longformer. Best fusion results per domain are in bold.
Target Training Data F1 Recall Precision vs Baseline
TAB TAB only (baseline) 0.859 0.937 0.793
TAB + Wiki (TW) 0.841 0.946 0.757 -2.1%
TAB + I2B2 (TI) 0.841 0.945 0.757 -2.1%
TAB + Wiki + I2B2 (TWI) 0.853 0.952 0.772 -0.7%
Wiki Wiki only (baseline) 0.828 0.986 0.713
Wiki + TAB (TW) 0.820 0.989 0.700 -1.0%
Wiki + I2B2 (WI) 0.842 0.985 0.735 +1.7%
Wiki + TAB + I2B2 (TWI) 0.849 0.973 0.753 +2.5%
I2B2 I2B2 only (baseline) 0.972 0.994 0.950
I2B2 + TAB (TI) 0.968 0.994 0.944 -0.4%
I2B2 + Wiki (WI) 0.970 0.993 0.948 -0.2%
I2B2 + TAB + Wiki (TWI) 0.967 0.991 0.944 -0.5%
Domain Characteristics Best Strategy Recall Effect Precision Effect Recommendation
TAB (Legal) All domains (TWI) +1.6% -2.6% Fusion beneficial
Wiki (Biographical) +Medical data (WI) -0.1% +3.1% Selective fusion
I2B2 (Medical) Domain-specific only Avoid fusion
  • Fusion combinations: TW (TAB+Wikipedia), TI (TAB+I2B2), WI (Wikipedia+I2B2), TWI (all three domains).
  • Results averaged across 7 random seeds. Positive percentages indicate improvement over single-domain baseline.
4.4 Few-Shot Learning Feasibility
Table 4 shows performance across varying training set sizes. TAB Domain exhibits non-monotonic relationships, achieving peak performance with medium-sized training sets ( 300 samples, -0.8% F1 vs. full data). Wikipedia Domain demonstrates optimal performance at  50% of full training data (200 samples, +2.1% F1), suggesting additional data may introduce complexity. I2B2 Domain shows strong positive correlation with training data volume, requiring substantial supervision due to domain complexity. Even with minimal training data (50 samples), models maintain reasonable performance: TAB F1 drops only 2.9%, Wikipedia decreases 1.6%, and I2B2 reduces 6.6%. These results validate few-shot PII recognition feasibility, particularly for less specialized domains.
Table 4: Few-Shot Learning Performance using Longformer. Results show F1/Recall/Precision for token-level evaluation across varying training set sizes. Best performance per domain is in bold.
Domain Training Size F1 Recall Precision vs Full Data Samples/Entity
TAB 50 samples 0.834 0.940 0.770 -2.9% 0.03
300 samples 0.852 0.947 0.774 -0.8% 0.19
550 samples 0.849 0.945 0.776 -1.2% 0.35
800 samples 0.848 0.934 0.775 -1.3% 0.52
1014 samples (full) 0.859 0.937 0.793 0.65
Wiki 50 samples 0.815 0.975 0.702 -1.6% 0.47
100 samples 0.819 0.979 0.704 -1.1% 0.93
200 samples 0.845 0.972 0.754 +2.1% 1.87
300 samples 0.836 0.982 0.728 +1.0% 2.80
453 samples (full) 0.828 0.986 0.713 4.23
I2B2 50 samples 0.908 0.982 0.844 -6.6% 0.17
100 samples 0.939 0.983 0.898 -3.4% 0.35
300 samples 0.966 0.991 0.942 -0.6% 1.04
500 samples 0.962 0.994 0.932 -1.0% 1.73
790 samples (full) 0.972 0.994 0.950 2.74
Domain Optimal Size Performance Pattern Min. Viable Size Data Sensitivity
TAB  30% (300 samples) Non-monotonic 50 samples Low
Wikipedia  44% (200 samples) Peak at mid-scale 50 samples Very Low
I2B2 100% (790 samples) Monotonic increase 300 samples High
Training Data Reduction TAB Wikipedia I2B2
90% reduction (50 samples) -2.9% F1 -1.6% F1 -6.6% F1
70% reduction (300/200/300) -0.8% F1 +2.1% F1 -0.6% F1
50% reduction (550/200/500) -1.2% F1 +2.1% F1 -1.0% F1
Few-Shot Feasibility Excellent Excellent Good
  • Results averaged across 7 random seeds. Samples/Entity shows training examples per entity type for context.
  • Optimal size represents training data volume that achieves best performance for each domain.
  • Minimum viable size indicates smallest dataset maintaining reasonable performance (>90% of full data F1).
4.5 Summary of Key Findings
Our comprehensive evaluation yields several critical insights:
  1. 1.
    Model Selection: Domain-tuned transformers, particularly Longformer, provide optimal privacy-utility balance;
  2. 2.
[…]
  3. 3.
    Data Fusion Strategy: Benefits vary by target domain—legal documents benefit from any additional data, biographical texts require strategic selection, medical records need domain-specific training;
  4. 4.
    Sample Efficiency: PII recognition demonstrates surprising robustness to data limitations, with acceptable performance using significantly reduced datasets.
5 Conclusions
This paper presents the first large-scale empirical study of cross-domain PII recognition, evaluating model transferability, data fusion strategies, and sample efficiency across healthcare, legal, and biographical texts. Domain-tuned transformers, particularly Longformer, consistently outperform rule-based and general NER models. Cross-domain transfer is highly asymmetric: legal-domain models transfer well to biographical data, but all models perform poorly on medical text. Multi-domain fusion offers gains for legal and biographical domains when done selectively but consistently harms performance on medical data. Effective PII recognition is also achievable with as little as 10% of training data in non-specialized domains.
These findings suggest legal-domain models may be repurposed for biographical texts with minimal loss, while medical applications require domain-specific training. While results are strong across many settings, future work should address limitations in handling complex linguistic structures and edge-case entities, extend evaluation to conversational and social media data, improve fusion strategies, and explore synthetic data for privacy-preserving few-shot learning.